In how many ways can you walk up a stairway that has 7 steps if you can take 1 or 2 steps at a time.
Answer by Theo(13342)

The total number of possible ways can be found as follows:

Find the possible combinations you can use to get to the number of steps.
Fin the possible number of ways each combination can be done.

Example:

Number of steps = 3
This can be done using the following possible combinations.
111
21

Add each number in each string and you get a total of 3.

111 can be done in 1 way.

21 can be done in 2 ways.

They would be 12 and 21.

Total number of possible ways to get up a flight of 3 steps is 3.

Those ways are:

111
12
21

Go to a flight of 4 steps.

Total number of possible combinations involving either 1 step or 2 steps are:

1111
22
211

you can do 1111 only 1 way.

you can do 22 only 1 way.

You can do 211 in 3 ways.

Those ways are:

211
121
112

There is therefore a total of 5 ways to get up a flight of 4 steps.

Those ways are

1111
22
211
121
112

Go to a flight of 5 steps.

Total number of possible combinations involving either 1 step or 2 steps are:

11111
1112
122

11111 can be done in only 1 way.

1112 can be done in 4 ways.

Those ways are:

1112
1121
1211
2111

122 can be done in 3 ways.

Those ways are:

122
212
221

Total number of ways to get up a flight of 5 steps is equal to:

1 + 4 + 3 = 8 ways

Those ways are:

11111
1112
1121
1211
2111
122
212
221

We could do 6, but you should have the idea by now, so we'll jump to 7 steps.

Possible combinations using only 1 step and 2 steps are:

1111111
111112
11122
1222

If you sum the digits in each string, you will get a total of 7 for each.

1111111 can be done in only 1 way.

111112 can be done in 6 ways.

Those ways are:

111112
111121
111211
112111
121111
211111

11122 can be done in 10 ways.

Those ways are:

11122
11212
11221
12112
12121
12211
21112
21121
21211
22111

1222 can be done in 4 ways.

those ways are:

1222
2122
2212
2221

The total number of ways you can get up a flight of 7 steps if you take 1 step at a time or 2 steps at a time would therefore be equal to:

1 + 6 + 10 + 4 = 21 ways.

There might be an easier way to do this but I haven't figured it out.

This way works, even though it does involve a little work.

How you determine the number of ways for each of the combinations is as follows:

Suppose one of the possible combinations for getting up a flight of 7 steps is as follows:

11122

That 3 of your strides that take 1 step at a time, and 2 of your strides that take 2 steps at a time.

You have taken a total of 5 strides.

The formula to determine how many possible ways you can do this is as follows:

Let n = total number of strides.
Let x = number of strides that take 2 steps at a time.
Let n-x = number of strides that take 1 step at a time.

The formula is n! / (x! * (n-x)!)

You should recognize that formula as the number of combinations formula (order is not important).

In this instance:

n = total number of strides = 5
x = total number of strides taking 2 steps at a time = 2
(n-x) = total number of strides taking 1 step at a time = 3.

The formula becomes:

5!/(2! * 3!)

This translates to 5*4*3! / (2*1*3!)

The 3! cancels out and you get:

5*4 / 2*1

This simplifies to 5*2 / 1 which equals 10.

That's how we got 10 possible ways to get the combination of 2 strides taking 2 steps at a time and 3 strides taking 1 step at a time which was shown as:

11122

We then determined what those possible ways could be.

This can be done with smaller numbers. It's much more difficult to do with larger number without the aid of a computer.

Those possible ways were:

11122
11212
11221
12112
12121
12211
21112
21121
21211
22111